4 - y = 14y find y
Answer by checkley79(3341)

4 - y = 14y
-Y-14Y=-4
-15Y=-4
Y=-4/-15
Y=4/15 ANS.
PROOF:
4-(4/15)=14(4/15)
(4*15-4)/15=(14*4)/15
(60-4)/15=56/15
56/15=56/15